# Sliced-Wasserstein Estimation with Spherical Harmonics as Control Variates

### v1

@@ Appendix A Algorithms for Sliced-Wasserstein Estimates
1:  Sample random projections $\theta_{1},\ldots,\theta_{n}\sim\mathcal{U}(\mathbb{S}^{d-1})$
@@
1:  Sample random projections $\theta_{1},\ldots,\theta_{n}\sim\mathcal{U}(\mathbb{S}^{d-1})$
@@
4:  Return ${\color[rgb]{0.80078125,0.46875,0.73828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.80078125,0.46875,0.73828125}\pgfsys@color@rgb@stroke{0.80078125}{0.46875}{0.73828125}\pgfsys@color@rgb@fill{0.80078125}{0.46875}{0.73828125}\mathrm{SHCV}}_{n}=\mathrm{I}_{n}^{\mathrm{ols}}$
@@
1:  Sample random projections $\theta_{1},\ldots,\theta_{n}\sim\mathcal{U}(\mathbb{S}^{d-1})$
@@
3:  Compute ${\color[rgb]{0.00390625,0.44921875,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.44921875,0.69921875}\pgfsys@color@rgb@stroke{0.00390625}{0.44921875}{0.69921875}\pgfsys@color@rgb@fill{0.00390625}{0.44921875}{0.69921875}\text{MC}}_{n}=(\mathds{1}_{n}/n)^{\top}f_{n}$
@@
7:  Compute $\gamma_{\text{low}}=\frac{\langle f_{n}-{\color[rgb]{0.00390625,0.44921875,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.44921875,0.69921875}\pgfsys@color@rgb@stroke{0.00390625}{0.44921875}{0.69921875}\pgfsys@color@rgb@fill{0.00390625}{0.44921875}{0.69921875}\text{MC}}_{n},C-b\rangle}{\|C-b\|_{2}^{2}}$ and control variate $\mathcal{C}=(\mathds{1}_{n}^{\top}(C-b))/n$
8:  Return ${\color[rgb]{0.0078125,0.62109375,0.44921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.0078125,0.62109375,0.44921875}\pgfsys@color@rgb@stroke{0.0078125}{0.62109375}{0.44921875}\pgfsys@color@rgb@fill{0.0078125}{0.62109375}{0.44921875}\text{CV}_{\text{low}}}={\color[rgb]{0.00390625,0.44921875,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.44921875,0.69921875}\pgfsys@color@rgb@stroke{0.00390625}{0.44921875}{0.69921875}\pgfsys@color@rgb@fill{0.00390625}{0.44921875}{0.69921875}\text{MC}}_{n}-\gamma_{\text{low}}\mathcal{C}$
@@
1:  Sample random projections $\theta_{1},\ldots,\theta_{n}\sim\mathcal{U}(\mathbb{S}^{d-1})$
@@
3:  Compute ${\color[rgb]{0.00390625,0.44921875,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.44921875,0.69921875}\pgfsys@color@rgb@stroke{0.00390625}{0.44921875}{0.69921875}\pgfsys@color@rgb@fill{0.00390625}{0.44921875}{0.69921875}\text{MC}}_{n}=(\mathds{1}_{n}/n)^{\top}f_{n}$
@@
7:  Compute $\gamma_{\text{up}}=\frac{\langle f_{n}-{\color[rgb]{0.00390625,0.44921875,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.44921875,0.69921875}\pgfsys@color@rgb@stroke{0.00390625}{0.44921875}{0.69921875}\pgfsys@color@rgb@fill{0.00390625}{0.44921875}{0.69921875}\text{MC}}_{n},C-b\rangle}{\|C-b\|_{2}^{2}}$ and control variate $\mathcal{C}=(\mathds{1}_{n}^{\top}(C-b))/n$
8:  Return ${\color[rgb]{0.87109375,0.5625,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.87109375,0.5625,0.01953125}\pgfsys@color@rgb@stroke{0.87109375}{0.5625}{0.01953125}\pgfsys@color@rgb@fill{0.87109375}{0.5625}{0.01953125}\text{CV}_{up}}={\color[rgb]{0.00390625,0.44921875,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.44921875,0.69921875}\pgfsys@color@rgb@stroke{0.00390625}{0.44921875}{0.69921875}\pgfsys@color@rgb@fill{0.00390625}{0.44921875}{0.69921875}\text{MC}}_{n}-\gamma_{\text{up}}\mathcal{C}$
Algorithm 7 Control Neighbors Sliced Wasserstein (Leluc et al., 2023)
@@
3:  Compute ${\color[rgb]{0.00390625,0.44921875,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.44921875,0.69921875}\pgfsys@color@rgb@stroke{0.00390625}{0.44921875}{0.69921875}\pgfsys@color@rgb@fill{0.00390625}{0.44921875}{0.69921875}\text{MC}}_{n}=(\mathds{1}_{n}/n)^{\top}f_{n}$
@@
5:  Return ${\color[rgb]{0.8359375,0.3671875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8359375,0.3671875,0}\pgfsys@color@rgb@stroke{0.8359375}{0.3671875}{0}\pgfsys@color@rgb@fill{0.8359375}{0.3671875}{0}\text{CVNN}}_{n}={\color[rgb]{0.00390625,0.44921875,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.44921875,0.69921875}\pgfsys@color@rgb@stroke{0.00390625}{0.44921875}{0.69921875}\pgfsys@color@rgb@fill{0.00390625}{0.44921875}{0.69921875}\text{MC}}_{n}-[\mathds{1}_{n}^{\top}\{\hat{f}_{n}-\mathrm{I}(\hat{f}_{n})\}]/n$
@@
6:  Return ${\color[rgb]{0.79296875,0.5703125,0.37890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.79296875,0.5703125,0.37890625}\pgfsys@color@rgb@stroke{0.79296875}{0.5703125}{0.37890625}\pgfsys@color@rgb@fill{0.79296875}{0.5703125}{0.37890625}\text{RQMC}}_{n}=(\mathds{1}_{n}/n)^{\top}f_{n}$
@@
• In Algorithm 5, with the notation $z=\bar{x}-\bar{y}$, the control variate is the centered version of the function $\theta\mapsto|\theta^{\top}z|^{2}$, which has known integral $\int_{\mathbb{S}^{d-1}}|\theta^{\top}z|^{2}\mathrm{d}\mathrm{P}(\theta)=\|z\|^{2}/d=b$ (Step 6 in the algorithm) and where $\mathrm{P}$ denotes the uniform distribution on $\mathbb{S}^{d-1}$. The Lower-CV estimate is thus an OLSMC estimate with control variate $\varphi(\theta)=|\theta^{\top}z|^{2}-\|z\|^{2}/d$, a polynomial in $\theta$ of degree two.
• Algorithm 6 is actually an OLSMC estimate with a single control variate given by the centered version of $\theta\mapsto\theta^{\top}M\theta$ where $M=M_{0}+M_{X}+M_{Y}$ with $M_{0}=(\bar{x}-\bar{y})(\bar{x}-\bar{y})^{\top}$, $M_{X}=\sum_{i=1}^{m_{X}}\alpha_{i}(x_{i}-\bar{x})(x_{i}-\bar{x})^{\top}$ and similarly for $M_{Y}$. The integral with respect to the uniform distribution on the sphere is $\int_{\mathbb{S}^{d-1}}\theta^{\top}M\theta\,\mathrm{d}\mathrm{P}(\theta)=\operatorname{Tr}(M)/d=b$ in Step 6 of the algorithm. The control variate is thus the centered quadratic polynomial $\varphi(\theta)=\theta^{\top}M\theta-\operatorname{Tr}(M)/d$.
• In Step 4 of Algorithm 7, $\hat{f}_{n}^{(i)}(\theta_{i})$ is a leave-one-out nearest-neighbor estimate computed with the Euclidean metric. In Step 5, computing $\mathrm{I}(\hat{f}_{n})$ requires sampling $M$ additional particles $\theta_{i}$ from $\mathrm{P}$ and we set $M=n^{1+2/d}$ as recommended by the theory in Leluc et al. (2023).
@@ Appendix B Numerical Details and Additional Results
To assess the finite-sample performance of the SHCV estimate, we first present some synthetic data examples involving the computation of $\operatorname{SW}_{2}^{2}(\mu,\nu)$ between multivariate Gaussian distributions. We consider different dimensions $d\in\{3;5;6;10;20\}$ and a varying number of random projections $n\in[10^{2};10^{4}]$. We consider the two cases of exact distributions in Section B.1 and sampled distributions in Section B.2. Then in Section B.3, we focus on the Sliced-Wasserstein distance $\operatorname{SW}_{2}(\mu_{m},\nu_{m})$ between 3D point clouds from the ShapeNetCore dataset of Chang et al. (2015). In both cases, the true value of the Sliced-Wasserstein distance is estimated with a very large number $n=10^{8}$ of particles. The different results report the mean squared error $\mathbb{E}[|{\operatorname{SW}}_{2,n}^{2}(\mu,\nu)-\operatorname{SW}_{2}^{2}(\mu,\nu)|^{2}]$ for the different estimates ${\operatorname{SW}}_{2,n}^{2}(\mu,\nu)$ where the expectation is computed as an average over $100$ independent runs. The experiments were performed on a laptop Intel Core i7-10510U CPU 1.80GHz $\times$ 8.
@@ B.1 Multivariate Gaussian: Exact Case with Distributions $\mu,\nu$ and Smooth Integrands
MC $2.6$e-$2$ $06.5\pm 2.3$ $8.4$e-$3$ $06.7\pm 0.7$ $2.0$e-$2$ $06.5\pm 0.3$
$\text{CV}_{low}$ $4.7$e-$3$ $18.1\pm 4.5$ $3.0$e-$3$ $18.4\pm 4.3$ $4.4$e-$3$ $18.9\pm 3.3$
$\text{CV}_{up}$ $1.5$e-$2$ $20.1\pm 5.2$ $7.0$e-$3$ $16.8\pm 3.2$ $1.6$e-$2$ $14.6\pm 1.6$
CVNN $7.7$e-$3$ $45.4\pm 12\phantom{.}$ $1.0$e-$2$ $62.0\pm 14\phantom{.}$ $2.6$e-$2$ $73.6\pm 12\phantom{.}$
QMC $4.9$e-$3$ $22.0\pm 7.1$ $4.1$e-$3$ $42.3\pm 9.7$ $1.3$e-$2$ $62.2\pm 6.1$
RQMC $7.5$e-$4$ $23.6\pm 6.7$ $2.4$e-$3$ $52.4\pm 12\phantom{.}$ $1.4$e-$2$ $84.3\pm 14\phantom{.}$
SHCV $\mathbf{4.8}$e-$\mathbf{5}$ $15.6\pm 4.5$ $\mathbf{1.7}$e-$\mathbf{3}$ $10.4\pm 4.9$ $\mathbf{2.7}$e-$\mathbf{3}$ $11.2\pm 5.2$
@@ B.2 Multivariate Gaussian: Sampled Case with Discrete Distributions $\mu_{m},\nu_{m}$
Table 4: Mean Squared Error (MSE) for sampled Gaussian distributions supported on $m=1000$ points in dimension $d\in\{5;10;20\}$ with $n\in\{500;1000\}$ particles, obtained over $100$ independent runs.
@@
Table 5: Computing time (ms) for sampled Gaussian distributions supported on $m=1000$ points in dimension $d\in\{5;10;20\}$ with $n\in\{500;1000\}$ particles, obtained over $100$ independent runs.
@@ B.3 Empirical Distributions of $3$D Point Clouds
We use three point clouds taken at random from the ShapeNetCore dataset of Chang et al. (2015) corresponding to the objects plane, lamp and bed, each composed of $m=2048$ points of $\mathbb{R}^{3}$ (see Figure 6 below for visualization). The SHCV estimate uses spherical harmonics up to degree $162L=$ 16 leading to $s=152$ control variates. Figure 7 below reports the boxplots of the error distribution $(\widehat{\operatorname{SW}}_{n}(\mu_{m},\nu_{m})-\operatorname{SW}(\mu_{m},\nu_{m}))$ for the different $\operatorname{SW}$ estimates based on $n$ random projections with $n\in\{100;250;500;1000\}$ obtained over $100$ independent runs while Table 6 presents the mean squared errors $\mathbb{E}[|\widehat{\operatorname{SW}}_{n}(\mu_{m},\nu_{m})-\operatorname{SW}(\mu_{m},\nu_{m})|^{2}]$ for the different $\operatorname{SW}$ estimates with $n\in\{100;500;1000\}$. Once again, our SHCV estimate provides the best performance with huge gains in terms of variance reduction.
@@
(a) plane/lamp
(b) lamp/bed
(c) plane/bed
Figure 7: Boxplots of the error distribution $(\widehat{\operatorname{SW}}_{n}(\mu_{m},\nu_{m})-\operatorname{SW}(\mu_{m},\nu_{m}))$ for different SW estimates based on $n$ random projections with $n\in\{100;250;500;1000\}$ obtained over $100$ independent runs.

### v2

@@ Acknowledgments
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Algorithms for Sliced-Wasserstein Estimates
1:  Sample random projections $\theta_{1},\ldots,\theta_{n}$ uniformly on $\mathbb{S}^{d-1}$
@@
1:  Sample random projections $\theta_{1},\ldots,\theta_{n}$ uniformly on $\mathbb{S}^{d-1}$
@@
4:  Return ${\color[rgb]{0.80078125,0.46875,0.73828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.80078125,0.46875,0.73828125}\mathrm{SHCV}}_{n}=\mathrm{I}_{n}^{\mathrm{ols}}$
@@
1:  Sample random projections $\theta_{1},\ldots,\theta_{n}$ uniformly on $\mathbb{S}^{d-1}$
@@
3:  Compute ${\color[rgb]{0.00390625,0.44921875,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.44921875,0.69921875}\text{MC}}_{n}=(\mathds{1}_{n}/n)^{\top}f_{n}$
@@
7:  Compute $\gamma_{\text{low}}=\frac{\langle f_{n}-{\color[rgb]{0.00390625,0.44921875,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.44921875,0.69921875}\text{MC}}_{n},C-b\rangle}{\|C-b\|_{2}^{2}}$ and control variate $\mathcal{C}=(\mathds{1}_{n}^{\top}(C-b))/n$
8:  Return ${\color[rgb]{0.0078125,0.62109375,0.44921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.0078125,0.62109375,0.44921875}\text{CV}_{\text{low}}}={\color[rgb]{0.00390625,0.44921875,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.44921875,0.69921875}\text{MC}}_{n}-\gamma_{\text{low}}\mathcal{C}$
@@
1:  Sample random projections $\theta_{1},\ldots,\theta_{n}$ uniformly on $\mathbb{S}^{d-1}$
@@
3:  Compute ${\color[rgb]{0.00390625,0.44921875,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.44921875,0.69921875}\text{MC}}_{n}=(\mathds{1}_{n}/n)^{\top}f_{n}$
@@
7:  Compute $\gamma_{\text{up}}=\frac{\langle f_{n}-{\color[rgb]{0.00390625,0.44921875,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.44921875,0.69921875}\text{MC}}_{n},C-b\rangle}{\|C-b\|_{2}^{2}}$ and control variate $\mathcal{C}=(\mathds{1}_{n}^{\top}(C-b))/n$
8:  Return ${\color[rgb]{0.87109375,0.5625,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.87109375,0.5625,0.01953125}\text{CV}_{up}}={\color[rgb]{0.00390625,0.44921875,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.44921875,0.69921875}\text{MC}}_{n}-\gamma_{\text{up}}\mathcal{C}$
Algorithm 7 Control Neighbors Sliced Wasserstein (Leluc et al., 2024)
@@
3:  Compute ${\color[rgb]{0.00390625,0.44921875,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.44921875,0.69921875}\text{MC}}_{n}=(\mathds{1}_{n}/n)^{\top}f_{n}$
@@
5:  Return ${\color[rgb]{0.8359375,0.3671875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8359375,0.3671875,0}\text{CVNN}}_{n}={\color[rgb]{0.00390625,0.44921875,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.44921875,0.69921875}\text{MC}}_{n}-[\mathds{1}_{n}^{\top}\{\hat{f}_{n}-\mathrm{I}(\hat{f}_{n})\}]/n$
@@
6:  Return ${\color[rgb]{0.79296875,0.5703125,0.37890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.79296875,0.5703125,0.37890625}\text{RQMC}}_{n}=(\mathds{1}_{n}/n)^{\top}f_{n}$
@@
• In Algorithm 5, with the notation $z=\bar{x}-\bar{y}$, the control variate is the centered version of the function $\theta\mapsto|\theta^{\top}z|^{2}$, which has known integral $\int_{\mathbb{S}^{d-1}}|\theta^{\top}z|^{2}\mathrm{d}\mathrm{U}(\theta)=\|z\|^{2}/d=b$ (Step 6 in the algorithm) and where $\mathrm{U}$ denotes the uniform distribution on $\mathbb{S}^{d-1}$. The Lower-CV estimate is thus an OLSMC estimate with control variate $\varphi(\theta)=|\theta^{\top}z|^{2}-\|z\|^{2}/d$, a polynomial in $\theta$ of degree two.
• Algorithm 6 is actually an OLSMC estimate with a single control variate given by the centered version of $\theta\mapsto\theta^{\top}M\theta$ where $M=M_{0}+M_{X}+M_{Y}$ with $M_{0}=(\bar{x}-\bar{y})(\bar{x}-\bar{y})^{\top}$, $M_{X}=\sum_{i=1}^{m_{X}}\alpha_{i}(x_{i}-\bar{x})(x_{i}-\bar{x})^{\top}$ and similarly for $M_{Y}$. The integral with respect to the uniform distribution on the sphere is $\int_{\mathbb{S}^{d-1}}\theta^{\top}M\theta\,\mathrm{d}\mathrm{U}(\theta)=\operatorname{Tr}(M)/d=b$ in Step 6 of the algorithm. The control variate is thus the centered quadratic polynomial $\varphi(\theta)=\theta^{\top}M\theta-\operatorname{Tr}(M)/d$.
• In Step 4 of Algorithm 7, $\hat{f}_{n}^{(i)}(\theta_{i})$ is a leave-one-out nearest-neighbor estimate computed with the Euclidean metric. In Step 5, computing $\mathrm{I}(\hat{f}_{n})$ requires sampling $M$ additional directions $\theta_{i}$ from $\mathrm{P}$ and we set $M=n^{1+2/d}$ as recommended by the theory in Leluc et al. (2024).
@@ Appendix B Numerical Details and Additional Results
To assess the finite-sample performance of the SHCV estimate, we first present some synthetic data examples involving the computation of $\operatorname{SW}_{2}^{2}(\mu,\nu)$ between multivariate Gaussian distributions. We consider different dimensions $d\in\{3;5;6;10;20\}$ and a varying number of random projections $n\in[10^{2};10^{4}]$. We consider the two cases of exact distributions in Section B.1 and sampled distributions in Section B.2. Then in Section B.3, we focus on the Sliced-Wasserstein distance $\operatorname{SW}_{2}(\mu_{m},\nu_{m})$ between 3D point clouds from the ShapeNetCore dataset of Chang et al. (2015). In both cases, the true value of the Sliced-Wasserstein distance is estimated with a very large number $n=10^{8}$ of projections. The different results report the mean squared error $\mathbb{E}[|{\operatorname{SW}}_{2,n}^{2}(\mu,\nu)-\operatorname{SW}_{2}^{2}(\mu,\nu)|^{2}]$ for the different estimates $\operatorname{SW}_{2,n}^{2}(\mu,\nu)$ where the expectation is computed as an average over $100$ independent runs. The experiments were performed on a laptop Intel Core i7-10510U CPU 1.80GHz $\times$ 8.
@@ B.1 Multivariate Gaussian: Exact Case with Distributions $\mu,\nu$ and Smooth Integrands
MC $2.6$e-$2$ $06.4\pm 1.1$ $8.4$e-$3$ $06.7\pm 0.8$ $2.0$e-$2$ $06.6\pm 0.5$
$\text{CV}_{low}$ $4.7$e-$3$ $11.1\pm 3.4$ $3.0$e-$3$ $11.6\pm 3.0$ $4.4$e-$3$ $12.1\pm 2.6$
$\text{CV}_{up}$ $1.5$e-$2$ $11.4\pm 3.5$ $7.0$e-$3$ $11.8\pm 3.1$ $1.6$e-$2$ $13.2\pm 2.8$
CVNN $7.7$e-$3$ $43.4\pm 13\phantom{.}$ $1.0$e-$2$ $60.0\pm 16\phantom{.}$ $2.6$e-$2$ $70.6\pm 14\phantom{.}$
QMC $4.9$e-$3$ $22.4\pm 3.4$ $4.1$e-$3$ $41.1\pm 11\phantom{.}$ $1.3$e-$2$ $60.9\pm 7.1$
RQMC $7.5$e-$4$ $25.6\pm 6.8$ $2.4$e-$3$ $40.9\pm 11\phantom{.}$ $1.4$e-$2$ $66.4\pm 14\phantom{.}$
SHCV $\mathbf{4.8}$e-$\mathbf{5}$ $09.8\pm 5.1$ $\mathbf{1.7}$e-$\mathbf{3}$ $08.9\pm 1.2$ $\mathbf{2.7}$e-$\mathbf{3}$ $14.3\pm 2.0$
@@ B.2 Multivariate Gaussian: Sampled Case with Discrete Distributions $\mu_{m},\nu_{m}$
Table 4: Mean Squared Error (MSE) for sampled Gaussian distributions supported on $m=1000$ points in dimension $d\in\{5;10;20\}$ with $n\in\{500;1000\}$ projections, obtained over $100$ independent runs.
@@
Table 5: Computing time (ms) for sampled Gaussian distributions supported on $m=1000$ points in dimension $d\in\{5;10;20\}$ with $n\in\{500;1000\}$ projections, obtained over $100$ independent runs.
@@ B.3 Empirical Distributions of $3$D Point Clouds
We use three point clouds taken at random from the ShapeNetCore dataset of Chang et al. (2015) corresponding to the objects plane, lamp and bed, each composed of $m=2048$ points of $\mathbb{R}^{3}$ (see Figure 6 below for visualization). The SHCV estimate uses spherical harmonics up to degree $162L=$ 16 leading to $s=152$ control variates. Figure 3 reports the boxplots of the error distribution $(\widehat{\operatorname{SW}}_{n}(\mu_{m},\nu_{m})-\operatorname{SW}(\mu_{m},\nu_{m}))$ for the different $\operatorname{SW}$ estimates based on $n$ random projections with $n\in\{100;250;500;1000\}$ obtained over $100$ independent runs while Table 6 presents the mean squared errors $\mathbb{E}[|\widehat{\operatorname{SW}}_{n}(\mu_{m},\nu_{m})-\operatorname{SW}(\mu_{m},\nu_{m})|^{2}]$ for the different $\operatorname{SW}$ estimates with $n\in\{100;500;1000\}$. Once again, our SHCV estimate provides the best performance with huge gains in terms of variance reduction.
@@
Method plane/lamp lamp/bed plane/bed
$n=100$ $n=500$ $n=1000$ $n=100$ $n=500$ $n=1000$ $n=100$ $n=500$ $n=1000$
MC $36.4\pm 10$ $184\pm 3$ $309\pm 3$ $37.9\pm 10$ $184\pm 3$ $308\pm 3$ $35.6\pm 10$ $185\pm 3$ $308\pm 3$
$\text{CV}_{\text{low}}$ $36.8\pm 10$ $180\pm 1$ $301\pm 1$ $38.4\pm 10$ $181\pm 2$ $300\pm 2$ $36.1\pm 11$ $181\pm 1$ $300\pm 1$
$\text{CV}_{\text{up}}$ $38.1\pm 10$ $187\pm 2$ $316\pm 1$ $39.7\pm 10$ $187\pm 1$ $315\pm 1$ $37.2\pm 11$ $188\pm 1$ $314\pm 1$
CVNN $40.4\pm 11$ $220\pm 1$ $417\pm 1$ $42.0\pm 11$ $220\pm 1$ $414\pm 2$ $39.5\pm 11$ $221\pm 1$ $414\pm 1$
QMC $32.9\pm 12$ $194\pm 3$ $324\pm 1$ $27.9\pm 10$ $194\pm 2$ $327\pm 1$ $38.0\pm 13$ $195\pm 2$ $325\pm 1$
RQMC $40.6\pm 11$ $202\pm 2$ $346\pm 2$ $42.0\pm 10$ $203\pm 4$ $344\pm 2$ $39.3\pm 11$ $202\pm 1$ $343\pm 1$
SHCV $37.9\pm 10$ $185\pm 2$ $311\pm 1$ $39.7\pm 10$ $185\pm 1$ $310\pm 1$ $37.1\pm 11$ $186\pm 1$ $310\pm 1$
Table 7: Computing times (ms) for different SW estimates based on $n$ random projections with $n\in\{100;500;1000\}$ obtained over $100$ independent runs.